# On the cardinality of measures of maximal relative entropy for smooth skew products

### v2

@@ 3. Topological Markov shifts
$\mathscr{E}_{s}=\left\{(P,R)\to(P^{\prime},R^{\prime})\;\middle|\;\begin{aligned} &(P,R),(P^{\prime},R^{\prime})\in\mathscr{V}_{s},\ P^{\prime}\cap\theta^{-1}(P)\neq\emptyset\\ &\text{and }\Theta^{-1}(R^{\prime})\cap R\neq\emptyset\end{aligned}\right\}.$
Define $(\Sigma_{s},\sigma_{s}):=(\Sigma_{s}(\mathscr{G}_{s}),\sigma_{s})$ and consider the map
$\pi_{s}:((P_{i},R_{i}))_{i\in\mathbb{Z}}\in\Sigma_{s}\mapsto\bigcap_{n\in\mathbb{N}}\overline{\bigcap_{-n\leq i\leq n}\Theta^{-i}(R_{i})}\in\Omega\times M.$
Observe that $\pi_{s}((P_{i},R_{i})_{i\in\mathbb{Z}})=\pi((R_{i})_{i\in\mathbb{Z}}).$
We divide the remaining of the proof into five steps.
@@
Step 2.
We show that if $\mu$ is a $\chi$-hyperbolic $\Theta$-invariant measure such that $\mu(\mathcal{X})=1$ then $\mu(\pi_{s}(\Sigma_{s}))=\mu(\pi_{s}(\Sigma_{s}^{\#}))=1,$ which implies item (2).
Recall from Theorem 3.4 (2) that if $\mu(\mathcal{X})=1$, then $\mu(\pi(\Sigma))=\mu(\pi(\Sigma^{\#}))=1$ for each $\chi$-hyperbolic $\Theta$-invariant measure. Since $\Sigma_{s}^{\#}\subset\Sigma_{s}$, it is enough to show that for each $\underline{R}=(R_{i})_{i\in\mathbb{Z}}\in\Sigma^{\#}$ there exists $\underline{P}=(P_{i})_{i\in\mathbb{Z}}\in\Sigma_{b}$ such that $\underline{(P,R)}=\{(P_{i},R_{i})\}_{i\in\mathbb{Z}}\in\Sigma_{s}^{\#}.$ Indeed, under this assumption we obtain
$\mu(\pi_{s}(\Sigma_{s}))\geq\mu(\pi_{s}(\Sigma_{s}^{\#}))\geq\mu(\pi(\Sigma^{\#}))=1$
@@
Let $(\omega,x)\in\Omega\times M$ such that $\pi(\underline{R})=(\omega,x)$. From the definition of $\pi$ we obtain that
$(\omega,x)\in\overline{\bigcap_{-n\leq i\leq n}\Theta^{-i}(R_{i})}\ \text{for each }n\in\mathbb{N}.$
In particular, $\Theta^{i}(\omega,x)=(\theta^{i}\omega,T_{\omega}^{i}(x))\in\overline{R_{i}}$ for each $i\in\mathbb{N}.$ For each $i\in\mathbb{N}$, let $P_{i}\in\mathscr{P}$ such that $\theta^{i}\omega\in P_{i}.$ Therefore, we obtain that
$\overline{R_{i}}\cap(P_{i}\times M)\neq\emptyset\ \text{for each }i\in\mathbb{N}.$
In this way,
$\overline{R_{i}}\cap(P_{i}\times M)\neq\emptyset,\,\Theta^{-1}(R_{i+1})\cap R_{i}\neq\emptyset\ \text{and}\ \theta^{-1}(P_{i+1})\cap P_{i}\neq\emptyset\ \text{for every }i\in\mathbb{Z}.$
Therefore $\underline{(P,R)}\in\Sigma_{s}$ and $\pi_{s}(\underline{(P,R)})=\pi(\underline{R})$, in the following we ensure that $\underline{(P,R)}\in\Sigma_{s}^{\#}$.
Recall that $\underline{R}\in\Sigma^{\#}$. Since $\underline{R}\in\Sigma^{\#}$ there exist $R^{+},R^{-}\in\mathscr{R}$ and increasing sequences of natural numbers $\{i_{k}\}_{k\in\mathbb{N}},\{j_{k}\}_{\in\mathbb{N}}\subset\mathbb{N}$ such that
$R_{i_{k}}=R^{+}\ \text{and }R_{-j_{k}}=R^{-}\ \text{for each }k\in\mathbb{N}.$
Recall that $\underline{P}\in\Sigma_{b}$ and $\Sigma_{b}$ is a shift which admits finitely many symbols. Then, there exists $P^{+},P^{-}\in\mathscr{P}$ and subsequences $\{i_{k}^{\prime}\}_{k\in\mathbb{N}}\subset\{i_{k}\}_{k\in\mathbb{N}}$ and $\{j_{k}^{\prime}\}_{\in\mathbb{N}}\subset\{j_{k}\}_{\in\mathbb{N}}$ such that
$(P,R)_{i_{k}^{\prime}}=(P_{i_{k}^{\prime}},R_{i_{k}^{\prime}})=(P^{+},R^{+})\ \text{and }(P,R)_{j_{k}^{\prime}}=(P_{-j_{k}^{\prime}},R_{-j_{k}^{\prime}})=(P^{-},R^{-})\ \text{for each }k\in\mathbb{N}.$
Thus, $\underline{(P,R)}\in\Sigma^{\#}_{s}$ and $\pi_{s}(\underline{(P,R)})=\pi(\underline{R}).$ Which concludes the proof of Step 2.
Step 3.
@@
$\mathrm{proj}_{\Sigma}:\ \Sigma_{s}\to\Sigma,\quad\big((P_{i},R_{i})_{i\in\mathbb{Z}}\big)\mapsto(R_{i})_{i\in\mathbb{Z}},$
@@
Let $(\omega,x)=\pi(\underline{R})$. Recall that $\pi_{b}:\Sigma_{b}\to\Omega$ is a finite-to-one surjective map over $NW(\theta)$. Therefore $\pi_{b}^{-1}(\omega)=\{\underline{P}^{(1)},\ldots,\underline{P}^{(k)}\}$ for some $k\in\mathbb{N}$. To conclude the first part of Step 3, observe that:
$\#\{\mathrm{proj}_{\Sigma}^{-1}(\underline{R})\}$ $=\#\left\{\underline{(P,R)}\in\Sigma_{s};\underline{P}\in\Sigma_{b}\right\}$
$=\#\left\{\underline{(P,R)}\in\Sigma_{s};\theta^{i}\omega\in P_{i}\ \text{and }\overline{R_{i}}\cap(P_{i}\times M)\neq\emptyset\ \text{for each }i\in\mathbb{Z}\right\}$
$\leq\#\left\{\underline{P}\in\Sigma_{b};\pi_{b}(\underline{P})=\omega\right\}=\#\pi_{b}^{-1}(\omega)\leq k_{\mathscr{P}}.$
@@
Hence $\pi_{s}:\Sigma_{s}^{\#}\to\mathcal{X}$ is finite-to-one. Independently, since $\pi_{b}:\Sigma_{b}\to\Omega$ is obtained from a Bowen partition for $\theta$, it is finite-to-one on $NW(\theta)$. This proves item (3).
Step 4.
@@
$\mathrm{proj}_{\Sigma_{b}}:\{(P_{i},R_{i})\}_{i\in\mathbb{Z}}\in\Sigma_{s}\mapsto\{(P_{i})\}_{i\in\mathbb{Z}}\in\Sigma_{b},$
and set $\Sigma_{b}^{\prime}=\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s}).$ Observe that $\mathrm{proj}_{\Sigma_{b}}$ is a Lipschitz $1$-block factor map, in fact:
$\mathrm{dist}_{\Sigma_{b}}(\mathrm{proj}_{\Sigma_{b}}(\underline{(P,R}),\mathrm{proj}_{\Sigma_{b}}(\underline{(P^{\prime},R^{\prime}})))=\mathrm{dist}_{\Sigma_{b}}(\underline{P},\underline{P^{\prime}})\leq\mathrm{dist}_{\Sigma}(\underline{(P,R)},\underline{(P^{\prime},R^{\prime})}).$ (6)
Let $\underline{(P,R)}\in\Sigma_{s}$ and $\pi_{s}(\underline{(P,R)})=(\omega,x)$. Therefore,
$\overline{R_{i}}\cap(P_{i}\times M)\neq\emptyset\ \text{and }\theta^{i}(\omega)\in P_{i}\text{ for every }i\in\mathbb{N}.$
In particular, we obtain that $\pi_{b}(\underline{P})=\omega.$
Hence
@@
Observe that $\mathrm{proj}_{\Sigma_{b}}$ is clearly surjective if $\mathcal{X}=\Omega\times M$, therefore $\Sigma_{b}^{\prime}=\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s})=\Sigma_{b}$. If $\mathcal{X}$ is a Borel homoclinic class, observe that, by construction
$\Sigma_{b}^{\prime}=\Sigma(\mathscr{G}^{\prime}),$ (7)
where $\mathscr{G}^{\prime}=(\mathscr{V}^{\prime},\mathscr{E}^{\prime})$ with
$\mathscr{V}^{\prime}=\{P\in\mathscr{P};\text{there exists }R\in\mathscr{R}\ \text{such that }\ \overline{R}\cap(P\times M)\neq\emptyset\},$
and
$\mathscr{E}^{\prime}=\left\{P\to P^{\prime}\;\middle|\;\begin{aligned} &P,P^{\prime}\in\mathscr{V}^{\prime},\ \theta^{-1}(P^{\prime})\cap P\neq\emptyset\ \text{and there exist }R,R^{\prime}\in\mathscr{R},\ \text{such that}\\ &\overline{R}\cap(P\times M),\overline{R^{\prime}}\cap(P^{\prime}\cap M)\neq\emptyset\ \text{and }\Theta^{-1}(R^{\prime})\cap R\neq\emptyset\end{aligned}\right\}.$
In this way, $\Sigma_{b}^{\prime}$ is a subshift of finite type.
Step 5.
@@
Observe that to prove item $(5)$ it is enough to show the existence of $\Sigma_{s}^{\prime}\subset\Sigma_{s}$ such that:
@@
  • $\Sigma_{b}^{\prime}=\mathrm{proj}_{\Sigma_{b}}(\Sigma^{\prime}_{s})\subset\Sigma_{b}$ is a subshift of finite type.
Since $\mathcal{X}$ is a Borel homoclinic class, recall that $\Sigma$ is an irreducible topological Markov shift. Therefore, there exists $\underline{y}\in\Sigma$ such that $\Sigma=\overline{\{\sigma^{i}\left(\underline{y}\right)\}_{i\in\mathbb{Z}}}.$ Take $\underline{x}\in\Sigma_{s}$ such that $\mathrm{proj}_{\Sigma}(\underline{x})=\underline{y}$. Let $\mathscr{G}^{\prime}_{s}\subset\mathscr{G}_{s}$ the maximal irreducible component of $\mathscr{G}_{s}$ containing $x_{0}$ (the 0-th element of $\underline{x}$) and define
$\Sigma^{\prime}_{s}:=\Sigma(\mathscr{G}^{\prime}_{s})\subset\Sigma(\mathscr{G}_{s})=\Sigma_{s}.$
Observe that $\underline{x}\in\Sigma_{s}^{\prime}$ and therefore $\sigma^{i}_{s}(\underline{x})\in\Sigma_{s}^{\prime}$ for each $i\in\mathbb{Z}.$ It follows that

### v3

@@ 3. Topological Markov shifts
$\{(P^{\prime},R^{\prime})\in\mathscr{V}_{s};(P^{\prime},R^{\prime})\to(P,R)\in\mathscr{E}_{s}\}\leq\#\{R^{\prime}\in\mathscr{V}_{s};R^{\prime}\to R\in\mathscr{E}\}\cdot\#\mathscr{P}<\infty.$
Therefore, $\Sigma_{s}$ is a locally compact topological Markov shift.
Given $(P_{i},R_{i})_{i\in\mathbb{Z}}\in\Sigma_{s}$ we define
$\pi_{s}((P_{i},R_{i})_{i\in\mathbb{Z}}):=\bigcap_{n\in\mathbb{N}}\overline{\bigcap_{-n\leq i\leq n}\Theta^{-i}[R_{i}\cap(\operatorname{int}P_{i}\times M)]}.$
We divide the remainder of the proof into five steps.
@@
We show that if given $m,n\in\mathbb{Z}$ with $n>m$ and a path
$(P_{m},R_{m})\to(P_{m+1},R_{m+1})\to\ldots\to(P_{n},R_{n})$
in $\mathscr{G}_{s}$ then
$\bigcap_{i=m}^{n}\Theta^{-i}[R_{i}\cap(\operatorname{int}P_{i}\times M)]\neq\varnothing.$ (5)
In particular, we obtain that for each $(P_{i},R_{i})_{i\in\mathbb{Z}}\in\Sigma_{s}$
$\pi_{s}:\Sigma_{s}$ $\to\Omega\times M$
$(P_{i},R_{i})_{i\in\mathbb{Z}}$ $\mapsto\bigcap_{n\in\mathbb{N}}\overline{\bigcap_{-n\leq i\leq n}\Theta^{-i}\left[R_{i}\cap(\operatorname{int}P_{i}\times M)\right]}$
is well defined and $\pi_{s}((P_{i},R_{i})_{i\in\mathbb{Z}})=\pi((R_{i})_{i\in\mathbb{Z}}).$
We follow [Sarig2013, Lemma 12.1] and argue by induction on the length of the path. Fix
$(P_{m},R_{m})\ \to\ (P_{m+1},R_{m+1})\ \to\ \cdots\ \to\ (P_{n},R_{n})\ \to\ (P_{n+1},R_{n+1}).$
If $n=m+1$ (i.e. a single edge from $m$ to $m+1$), the claim regarding the nonemptiness of the intersection in (5) is immediate from the definition of the edge $(P_{m},R_{m})\to(P_{m+1},R_{m+1})$.
Assume the statement holds up to $n$, and prove it for $n+1$. By the induction hypothesis, there exists $(\omega_{0},x_{0})$ such that, writing $(\omega_{i},x_{i}):=\Theta^{i}(\omega_{0},x_{0})$,
$(\omega_{i},x_{i})\in R_{i}\cap\big(\operatorname{int}P_{i}\times M\big)\ \text{for all }i=m,\dots,n.$
Since $(P_{n},R_{n})\to(P_{n+1},R_{n+1})$, choose
$(\upsilon_{n},y_{n})\in\big(R_{n}\cap(\operatorname{int}P_{n}\times M)\big)\cap\Theta^{-1}\big(R_{n+1}\cap(\operatorname{int}P_{n+1}\times M)\big).$
Set $(\zeta_{n},z_{n}):=[\,(\upsilon_{n},y_{n}),(\omega_{n},x_{n})\,]_{\Theta}\in R_{n},$ which is well defined because $R_{n}$ is a rectangle with local product structure. Since $\Theta$ is a skew product and $(\upsilon_{n},y_{n}),(\omega_{n},x_{n})\in R_{n}\cap(\operatorname{int}P_{n}\times M)$, the bracket decouples in base and fibre:
$(\upsilon_{n},y_{n}),(\omega_{n},x_{n})\,]_{\Theta}=\big(\,[\upsilon_{n},\omega_{n}]_{\theta},\,z_{n}\big),$
and in particular $\zeta_{n},\omega_{n}]_{\theta}\in\operatorname{int}P_{n}$.
Finally, define $(\zeta_{0},z_{0}):=\Theta^{-n}(\zeta_{n},z_{n}).$ In the following, we show that
$(\zeta_{0},z_{0})\in\bigcap_{i=m}^{n+1}\Theta^{-i}\left[R_{i}\cap\big(\operatorname{int}P_{i}\times M\big)\right],$
which is sufficient to establish (5).
Define $(\zeta_{i},z_{i}):=\Theta^{i}(\zeta_{0},z_{0})$ for each $i\in\{m,\ldots,n\}$. For clarity in what follows, we denote by $W^{s}_{\Theta}$ (resp. $W^{u}_{\Theta}$) the local stable (resp. unstable) manifold of $\Theta$, and by $W^{s}_{\theta}$ (resp. $W^{u}_{\theta}$) the local stable (resp. unstable) manifold of $\theta$. Using the Markov properties of the partitions $\mathscr{R}$ and $\mathscr{P}$, we obtain
  • $\Theta^{n+1}(\zeta_{0},z_{0})=\Theta(\zeta_{n},z_{n})\in R_{n+1}\cap(\operatorname{int}P_{n+1}\times M)$, because
    $\Theta[W^{s}_{\Theta}((\zeta_{n},z_{n}),R_{n})]\subset W^{s}_{\Theta}(\Theta(\zeta_{n},z_{n}),R_{n+1})\subset R_{n+1}$
    and
    $\theta[W^{s}_{\theta}(\zeta_{n},\operatorname{int}P_{n})]\subset W^{s}_{\theta}(\theta(\zeta_{n}),\operatorname{int}P_{n+1})\subset\operatorname{int}P_{n+1}.$
  • $\Theta^{n}(\zeta_{0},z_{0})=(\zeta_{n},z_{n})\in R_{n}\cap(\operatorname{int}P_{n}\times M)$ by construction.
  • $\Theta^{n-1}(\zeta_{0},z_{0})=\Theta^{-1}(\zeta_{n},z_{n})\in R_{n-1}\cap(\operatorname{int}P_{n-1}\times M)$, because
    $\Theta^{n}(\zeta_{0},z_{0})\in W^{u}_{\Theta}((\zeta_{n},z_{n}),R_{n})\subset R_{n}$
    hence
    $\Theta^{n-1}(\zeta_{0},z_{0})\in\Theta^{-1}[W^{u}_{\Theta}((\zeta_{n},z_{n}),R_{n})]\subset W^{u}_{\Theta}((\zeta_{n-1},z_{n-1}),R_{n-1})\subset R_{n-1},$
    and on the base,
    $\theta^{n}(\zeta_{0})\in W^{u}_{\theta}(\zeta_{n},\operatorname{int}P_{n})\subset\operatorname{int}P_{n}$
    so
    $\theta^{n-1}(\zeta_{0})\in\theta^{-1}[W^{u}_{\theta}(\zeta_{n},\operatorname{int}P_{n})]\subset W^{u}_{\theta}(\zeta_{n-1},\operatorname{int}P_{n-1})\subset\operatorname{int}P_{n-1}.$
  • $\Theta^{n-2}(\zeta_{0},z_{0})\in R_{n-2}\cap(\operatorname{int}P_{n-2}\times M)$, because from the previous item
    $\Theta^{n-1}(\zeta_{0},z_{0})\in W^{u}((\zeta_{n-1},z_{n-1}),R_{n-1})\subset R_{n-1}$
    and therefore
    $\Theta^{n-2}(\zeta_{0},z_{0})\in\Theta^{-1}[W^{u}_{\Theta}((\zeta_{n-1},z_{n-1}),R_{n-1})]\subset W^{u}_{\Theta}((\zeta_{n-2},z_{n-2}),R_{n-2})\subset R_{n-2},$
    with the analogous inclusion on the base:
    $\theta^{n-2}(\zeta_{0})\in\theta^{-1}[W^{u}_{\theta}(\zeta_{n-1},\operatorname{int}P_{n-1})]\subset W^{u}_{\theta}(\zeta_{n-2},\operatorname{int}P_{n-2})\subset\operatorname{int}P_{n-2}.$
Continuing in this way, we obtain
$\Theta^{n+1-k}(\zeta_{0},z_{0})\in R_{n+1-k}\cap(\operatorname{int}P_{n+1-k}\times M)\text{ for all }0\leq k\leq n+1-m.$
which shows (5).
We now conclude Step 1. By (5) we have
$\emptyset\neq\pi_{s}(\{(P_{i},R_{i})\}_{i\in\mathbb{Z}})$ $=\bigcap_{n\in\mathbb{N}}\overline{\bigcap_{-n\leq i\leq n}\Theta^{-i}\left[R_{i}\cap(\operatorname{int}P_{i}\times M)\right]}$
$\subset\bigcap_{n\in\mathbb{N}}\overline{\bigcap_{-n\leq i\leq n}\Theta^{-i}(R_{i})}=\pi((R_{i})_{i\in\mathbb{Z}}).$
Since $\pi((R_{i})_{i\in\mathbb{Z}})$ is a singleton, it follows that $\pi_{s}(\{(P_{i},R_{i})\}_{i\in\mathbb{Z}})=\pi((R_{i})_{i\in\mathbb{Z}}).$
Step 2.
@@
Step 3.
We show that if $\mu$ is a $\chi$-hyperbolic, $\Theta$-invariant measure with $\mu(\mathcal{X})=1$ and $(\mathrm{proj}_{\Omega})_{*}\mu=\mathbb{P}$, then $\mu(\pi_{s}(\Sigma_{s}))=\mu(\pi_{s}(\Sigma_{s}^{\#}))=1$. Consequently, item (2) follows.
Recall from Theorem 3.5 (2) that if $\mu(\mathcal{X})=1$, then $\mu(\pi(\Sigma))=\mu(\pi(\Sigma^{\#}))=1$ for each $\chi$-hyperbolic $\Theta$-invariant measure. Since $\Sigma_{s}^{\#}\subset\Sigma_{s}$, and $(\mathrm{proj}_{\Omega})_{*}\mu=\mathbb{P}$ we have from (4) that
$\mu\left((\omega,x)\in\Omega\times M;\,\pi(\underline{R})=(\omega,x)\ \text{for some }\underline{R}\in\Sigma^{\#}\ \text{and }\omega\not\in\bigcup_{i\in\mathbb{Z}}\theta^{-i}(\partial\mathscr{P})\right)=1.$
Therefore, it is enough to show that for each
$\underline{R}=(R_{i})_{i\in\mathbb{Z}}\in\Sigma^{\#}\ \text{such that }\mathrm{proj}_{\Omega}\circ\pi(\underline{R})\not\in\bigcup_{i\in\mathbb{Z}}\theta^{-i}(\partial\mathscr{P}),$
there exists $\underline{P}=(P_{i})_{i\in\mathbb{Z}}\in\Sigma_{b}$ such that $\underline{(P,R)}=\{(P_{i},R_{i})\}_{i\in\mathbb{Z}}\in\Sigma_{s}^{\#}.$ Indeed, under this assumption, we obtain
$\mu(\pi_{s}(\Sigma_{s}))$ $\geq\mu(\pi_{s}(\Sigma_{s}^{\#}))$
$\geq$ $\mu\left((\omega,x)\in\Omega\times M;\,\pi(\underline{R})=(\omega,x)\ \text{for some }\underline{R}\in\Sigma^{\#}\ \text{and }\omega\not\in\bigcup_{i\in\mathbb{Z}}\theta^{-i}(\partial\mathscr{P})\right)$
$=$ $1.$
@@
Assume that $\underline{R}\in\Sigma^{\#}$ and $\pi(\underline{R})=(\omega,x)\in\left(\Omega\setminus\bigcup_{i\in\mathbb{Z}}\theta^{-i}(\partial\mathscr{P})\right)\times M$. For each $i\in\mathbb{Z}$ let $P_{i}\in\mathscr{P}$ be such that $\theta^{i}(\omega)\in\operatorname{int}P_{i}.$ Since $\theta:\Omega\to\Omega$ is Anosov we have that
$\bigcap_{-n\leq i\leq n}\theta^{-i}(\operatorname{int}P_{i})\times M=\bigcap_{-n\leq i\leq n}\Theta^{-i}(\operatorname{int}P_{i}\times M)$
is an open neighbourhood of $(\omega,x)$ in $\Omega\times M.$ Since $\pi(\underline{R})=(\omega,x)$ it follows that
$\bigcap_{-n\leq i\leq n}\Theta^{-i}(R_{i})\cap V\neq\varnothing$
for any open neighbourhood $V$ of $(\omega,x)$ in $\Omega\times M$. In particular,
$\varnothing\neq\bigcap_{-n\leq i\leq n}\Theta^{-i}(R_{i})\cap\bigcap_{-n\leq i\leq n}\Theta^{-i}\left(\operatorname{int}P_{i}\times M\right)=\bigcap_{-n\leq i\leq n}\Theta^{-i}\left[R_{i}\cap(\operatorname{int}P_{i}\times M)\right].$
Therefore, we obtain that $\{(P_{i},R_{i})\}_{i\in\mathbb{Z}}\in\Sigma_{s}.$ Since $\Sigma_{b}$ admits finitely many symbols and $\underline{R}\in\Sigma^{\#}$, by the pigeonhole principle, it is clear that $\{(P_{i},R_{i})\}_{i\in\mathbb{Z}}\in\Sigma_{s}^{\#}.$
Step 4.
@@
$\mathrm{proj}_{\Sigma}:\Sigma_{s}\to\Sigma,\ \big((P_{i},R_{i})_{i\in\mathbb{Z}}\big)\mapsto(R_{i})_{i\in\mathbb{Z}},$
@@
Recall that $\pi_{b}:\Sigma_{b}\to\Omega$ is a finite-to-one surjective map over $\Omega$. Given $\underline{(P,R)}=(P_{i},R_{i})_{i\in\mathbb{Z}}\in\Sigma_{s}$, on the one hand we have that
$\pi_{s}(\underline{(P,R)})=\bigcap_{n\in\mathbb{N}}\overline{\bigcap_{-n\leq i\leq n}\Theta^{-i}\left[R_{i}\cap(\operatorname{int}P_{i}\times M)\right]}\subset\bigcap_{i\in\mathbb{Z}}\theta^{-n}(P_{n})\times M=\pi_{b}(\underline{P})\times M.$
On the other hand, we obtain that from Step 1 we have that $\pi_{s}(\underline{P,R})=\pi(\underline{R}).$ It follows that
$\#\{\mathrm{proj}_{\Sigma}^{-1}(\underline{R})\}$ $\leq\#\left\{\underline{(P,R)}\in\Sigma_{s};\mathrm{proj}_{\Omega}\circ\pi(\underline{R})=\pi_{b}(\underline{P})\right\}$
$=\sup_{\omega\in\Omega}\#\pi^{-1}(\omega)\leq\kappa_{\mathscr{P}}<\infty$
@@
Hence $\pi_{s}:\Sigma_{s}^{\#}\to\mathcal{X}$ is finite-to-one. Independently, since $\pi_{b}:\Sigma_{b}\to\Omega$ is obtained from a Markov partition for $\theta$, it is finite-to-one on $\Omega$. This proves item (3).
Step 5.
We show that for any $\chi$-hyperbolic $\mu\in\mathcal{M}_{e}(\Theta)$ such that $(\mathrm{proj}_{\Omega})_{*}\mu=\mathbb{P}$, there exists $\widehat{\mu}\in M_{e}(\sigma_{s})$ such that $(\pi_{s})_{*}\widehat{\mu}=\mu$.
This is done exactly as in [Sarig2013, Proof of Proposition 13.1]; we sketch the argument. By [Sarig2013, Step 1 of Proposition 13.1] the map
$F:(\omega,x)\in\pi_{s}(\Sigma_{s}^{\#})\longmapsto\#\Big(\big.\pi_{s}^{-1}\big|_{\Sigma_{s}^{\#}}\{(\omega,x)\}\Big)\in\mathbb{N}$
is measurable. Moreover, there exists $k=k_{\mu}$ such that $F(\omega,x)=k$ for $\mu$-almost every $(\omega,x)\in\pi_{s}(\Sigma_{s}^{\#})$ ([Sarig2013, Step 1 of Proposition 13.1]; see also [Yoo2018, Lemma 3.1]). Define the measure on $\Sigma_{s}^{\#}$ by
$\widehat{\mu}(E):=\int_{\Omega\times M}\frac{1}{F(\omega,x)}\left(\sum_{\pi_{s}\underline{(P,R)}=(\omega,x)}\mathbbm{1}_{E}\big(\underline{(P,R)}\big)\right)\,\mathrm{d}\mu(\omega,x),$
for any Borel set $E\subset\Sigma_{s}^{\#}$. By [Sarig2013, Steps 3–4 of Proposition 13.1], the measure $\widehat{\mu}$ is
  • a well-defined $\sigma_{s}$-invariant probability measure; and
  • such that $(\pi_{s})_{*}\widehat{\mu}=\mu$ and $h_{\widehat{\mu}}(\sigma_{s})=h_{\mu}(\Theta)$.
Observe that, a priori, there is no need for $\widehat{\mu}$ to be ergodic. The proof is finished by observing that each (typical) $\sigma_{s}$-invariant ergodic component $\widehat{\mu}_{i}$ in the ergodic decomposition of $\widehat{\mu}$ also satisfies $(\pi_{s})_{*}\widehat{\mu}_{i}=\mu$ and $h_{\widehat{\mu}_{i}}(\sigma_{s})=h_{\mu}(\Theta)$ (see [Sarig2013, Step 5]).
Step 6.
@@
$\mathrm{proj}_{\Sigma_{b}}:\ \{(P_{i},R_{i})\}_{i\in\mathbb{Z}}\in\Sigma_{s}\ \mapsto\ \{P_{i}\}_{i\in\mathbb{Z}}\in\Sigma_{b}.$
Observe that $\mathrm{proj}_{\Sigma_{b}}$ is a Lipschitz $1$-block map; in fact:
$\mathrm{dist}_{\Sigma_{b}}\big(\mathrm{proj}_{\Sigma_{b}}(\underline{(P,R)}),\mathrm{proj}_{\Sigma_{b}}(\underline{(P^{\prime},R^{\prime})})\big)=\mathrm{dist}_{\Sigma_{b}}(\underline{P},\underline{P^{\prime}})\leq\mathrm{dist}_{\Sigma_{s}}(\underline{(P,R)},\underline{(P^{\prime},R^{\prime})}).$ (8)
Let $\underline{(P,R)}\in\Sigma_{s}$ and $\pi_{s}(\underline{(P,R)})=(\omega,x)$. Observe that, by construction, $\pi_{b}(\underline{P})=\omega$. Hence
@@
Observe that, by construction, $\mathrm{proj}_{\Sigma_{b}}\circ\sigma_{s}(\underline{(P,R)})=\sigma_{b}(\underline{P})=\theta\circ\mathrm{proj}_{\Sigma_{b}}(\underline{(P,R)})$.
Let $\mu$ be a $\chi$-hyperbolic measure on $\Omega\times M$ such that $(\mathrm{proj}_{\Omega})_{*}\mu=\mathbb{P}$. From Step 5, there exists $\widehat{\mu}\in\mathcal{M}_{e}(\sigma_{s})$ such that $(\pi_{s})_{*}\widehat{\mu}=\mu$. We therefore obtain that
$\mathbb{P}$ $=(\mathrm{proj}_{\Omega})_{*}\mu=(\mathrm{proj}_{\Omega})_{*}(\pi_{s})_{*}\widehat{\mu}=(\pi_{b})_{*}\big[(\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}\big$].
Defining the $\sigma_{b}$-invariant measure $\widehat{\mathbb{P}}:=(\mathrm{proj}_{\Sigma_{b}})_{*}\widehat{\mu}$, we have $(\pi_{b})_{*}\widehat{\mathbb{P}}$$\mathbb{P}$. Since $\mathbb{P}$ has full support, (4) ensures that $\mathbb{P}\big[\bigcup_{i\in\mathbb{Z}}\theta^{i}\partial\mathscr{P}\big]=0$, and therefore $\widehat{\mathbb{P}}$ is the unique $\sigma_{b}$-invariant lift of $\mathbb{P}$ with respect to the semi-conjugacy $\pi_{b}$ (see [BowenBook, Theorem 3.18]), which implies (4).
Step 7.
@@
Let $\mathcal{X}$ be a Borel homoclinic class. We restrict attention to the Borel homoclinic classes $\mathcal{X}$ with the property that there exists a $\chi$-hyperbolic $\Theta$-invariant ergodic probability measure $\mu$ supported on $\mathcal{X}$ such that $(\mathrm{proj}_{\Omega})_{*}\mu=\mathbb{P}$. Recall that $\mathrm{supp}\,\mathbb{P}=\Omega$. To prove item $(5)$ it suffices to find a subshift $\Sigma_{s}^{\prime}\subset\Sigma_{s}$ such that
@@
  • $\widehat{\mathbb{P}}[\mathrm{proj}_{\Sigma_{b}}(\Sigma^{\prime}_{s})]=1$, where $\widehat{\mathbb{P}}$ is the unique measure in $\mathcal{M}_{e}(\Sigma_{b})$ such that $(\pi_{b})_{*}\widehat{\mathbb{P}}$$\mathbb{P}$.
Since $\mathcal{X}$ is a Borel homoclinic class, we may (and do) take $\Sigma$ to be an irreducible two-sided topological Markov shift. Then $(\Sigma,\sigma)$ is topologically transitive and $\sigma$ is a homeomorphism. Assume $\Sigma$ is infinite (the periodic-orbit case is trivial).
Because $\Sigma$ is a two-sided countable Markov shift with the standard metric (see Definition 3.1), it is completely metrizable and has no isolated points. By transitivity, there exist $\underline{R^{0}}$ with dense forward orbit and $\underline{R^{1}}$ with dense backward orbit (see [SarigTF-TMS-2009, Proposition 1.1 and footnote 5]).
By [Manoussos, Thm. 2.4] applied to $\sigma:\Sigma\to\Sigma$ and to $\sigma^{-1}:\Sigma\to\Sigma$, we obtain that the sets